Two children, Angelica (mass 35 kg ) and Boris (mass 50 kg ), are playing on a merry-go-round (which you can assume is a solid disk with mass 200 kg  and radius 1.5m). Assume that any friction on the axle of the merry-go-round is negligible.

A.) Angelica starts spinning the merry-go-round, giving it an angular velocity of 8 rad/s , then she stops pushing it. Boris runs with a speed of 9 m/s directly toward the center of the merry-go-round. He jumps and lands on the edge of the merry-go-round. What is the angular speed of the merry-go-round once Boris lands on it?

B.) After landing, Boris walks to the center of the merry-go-round. What is the angular speed of the merry-go-round when he gets to the center?

C.) While Boris is in the center of the merry-go-round, Angelica runs up with a speed of 6 m/s  tangentially in the same direction the merry-go-round is spinning, jumping onto the edge of the merry-go-round. What is the angular speed of the merry-go-round after Angelica lands on it?

D.) Angelica stays in place while Boris walks back to the edge of the merry-go-round. What is the angular speed of the merry-go-round when he gets there?
Answer #1

a)

using conservation of angular momentum

0.5 M r^2 w1 = (0.5 M r^2 + m r^2) w'

0.5* 200* 8 = ( 0.5* 200 + 50) w'

w' = 5.33 rad/s

======

b)

w = 8 rad/s

======

c)

again using momentum conservation

I w1 + mvr = (I + mr^2) w'

0.5* 200* 1.5^2 * 8 + 35*6* 1.5 = ( 0.5* 200* 1.5^2 + 35* 1.5^2)w'

w' = 6.963 rad/s

d)

using conservation of angular momentum

(I + ma r^2) w' = ( I + (ma + mb) r^2) w

( 0.5* 200* 1.5^2 + 35* 1.5^2)*6.963 = (0.5* 200* 1.5^2 + ( 35 + 50)* 1.5^2)* wf

wf = 5.081 rad/s

======

Comment before rate in case any doubt, will reply for sure.. goodluck